Two cards are drawn at random from a set of 52 cards.What is the probability that the both cards are either kings or queens

1
Expert's answer
2022-04-18T16:44:19-0400

P(two kings or two queens) = p(2 kings ) + p(2 queens)

"p(2 queens) = (4/52) \cdot (3/51) = 0.0045"

"p(2queens)=(4\/52)\u22c5(3\/51)=0.0045"

"0.0045 + 0.0045 = 0.009"